OpenStudy (anonymous):

How to solve for t: 0.5=0.7^(t/15). I would appreciate it more if you help me how to do it, not give the exact answer right away :)

OpenStudy (anonymous):

apply log

OpenStudy (anonymous):

uhm, how? I'm very new

OpenStudy (anonymous):

log .5 = t/15 log .7?

OpenStudy (anonymous):

change of base

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

t= 15 ln (0.5) / ln 0.7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks kind sir

OpenStudy (anonymous):

yw